OpenStudy (dumbledore18):

Check my answers? Find the domain of y= 2x-4 (square root symbol is over 2x-4) My answer is X= -2 A square garden plot has an area of 24 ft2. Find the length of each side in simplest radical form. Calculate the length of each side to the nearest tenth of a foot. my answer is 24 over 4 ; 6 ft. Thanks

OpenStudy (photon336):

\[\frac{ 2x-4 }{ \sqrt{2x-4} } = f(x)\] right?

OpenStudy (photon336):

@Dumbledore18

OpenStudy (dumbledore18):

Yes.

OpenStudy (photon336):

look at the denominator, what do you know about radical signs?

OpenStudy (dumbledore18):

Pretty much nothing... Lol

OpenStudy (photon336):

lol well #1 thing you cant have a negative number under a radical sign, well you can but it won't be a real number, \[can't~have~any~number~less~than~zero~under~radical \sqrt{x<0}\]

OpenStudy (photon336):

so this means that \[\sqrt{2x-4} > 0\]

OpenStudy (marcelie):

https://www.youtube.com/watch?v=hZEGZMb4uzQ this might be useful to find domain as well

OpenStudy (photon336):

\[2x-4 > 0\] \[x > 2 \] so this means that any number greater than 2 the function will be defined. can you see why?